OpenStudy (anonymous):

In a right triangle, find the length of the side not given b=1 c=square root 17

OpenStudy (anonymous):

a^2+b^2=c^2 \[a ^{2}+1^{2}=\sqrt{17}^{2}\] \[a^{2}=17-1\] \[a=4\]

OpenStudy (amistre64):

that of course assumes that a, b, and c are determined by the convention of writing the pythag as a^2 + b^2 = c^2 :)

OpenStudy (amistre64):

in an ellipse, a is the hypotenuse and c is a leg ... by convention

OpenStudy (anonymous):

That thought did occur to me. I went with the assumption that since b<c that c would be the hypotenuse. It's possible that a was the hypotenuse :)

OpenStudy (anonymous):

That and the numbers played nicely :p